OpenStudy (samsan9):

Identify the shape of this system. { x^2+y^2+2x+12y=0 { x^2y^2+4x+6y+12=0

OpenStudy (anonymous):

polynomial?

OpenStudy (anonymous):

im just throwing something out there.

OpenStudy (samsan9):

i think its supposed to be a conic section

OpenStudy (mathmate):

Is there a + sign after the x^2 of the second equation, as follows: \(x^2+y^2+2x+12y=0\) \( x^2+y^2+4x+6y+12=0\) They are both ellipses centred not at origin. Standard Equation of ellipse centred at (h,k) \(\large \frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\)

OpenStudy (samsan9):

yes it has a + in the middle and so its an ellipse :o?

OpenStudy (mathmate):